
The elusive issue with the empty header for the spectra appeared again and affected 4 frames. Then it dissapeared without notice... This time I was able to get a snapshot of the DHS window showing an error, so that may be helpful (I can send those via email tomorrow). Other than that, throughout the night we occasionally ran into "highlighted hard points" for M2. The effects on the images and spectra were measurable but minimal, so I decided to keep observing instead of moving to zenith every time and do a reset.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2022-04-06 | 10.5 | 0.0 | 0.0 | 0.0 | 0.0 | 10.5 |
Links
[1] mailto:vinicius.placco@noirlab.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/40859/delete?destination=print/node/18793
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/40859/edit?destination=print/node/18793
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/18793?destination=print/node/18793
Besides the above, pretty uneventful night (which is good!) with very good weather.